Measures of liquidity, Solvency, and Profitability The comparative financial statements of Marshall Inc. are as follows. TheMarshall Inc. Comparative Income Statement For the Years Ended December 31, 20Y2 and 2041 20Y2 20Y1 Sales $4,135,450 $3,810,120Y2 20Y1 Assets Current assets Cash $660,400 $781,480 Marketable securities 999,520 1,295,020 730,000 Accounts receivable (n$680,000 Preferred $0.70 stock, $40 par Common stock, $10 par $680,000 780,000 780,000 3,892,600 3,187,400 Retained earnings2. Current ratio 3. Quick ratio 4. Accounts receivable turnover 5. Number of days sales in receivables days 6. Inventory tur
Answer #1
Marshall Inc.
Year 20Y2
1) Working Capital=Current Assets-Current Liabilities
Total Current Assets=(A) $                                                             31,42,656.00
Total Current Liabilities=(B) $                                                               9,82,080.00
Working Capital=(A)-(B) $                                                             21,60,576.00
2) Total Current Assets=(A) $                                                             31,42,656.00
Total Current Liabilities=(B) $                                                               9,82,080.00
Current Ratio=Current Assets/Current Liabilities=(A)/(B) 3.20
3) Quick Ratio=(Total Current Assets-Inventory-Prepaid Expenses)/Current Liabilities ($3142656-$584000-$124936)/$982080
2.478
4) Accounts Receivable Turnover=Credit Sales/Average Accounts Receivable
Credit Sales=(A) $                                                             41,35,450.00
Beginning Accounts Receivable $                                                               7,30,000.00
Ending Accounts Receivable $                                                               7,73,800.00
Average Accounts Receivable=(Beginning Accounts Receivable+Ending Accounts Receivable)/2=(B) $                                                               7,51,900.00
Accounts Receivable Turnover=(A)/(B) 5.50 Times
5) Number of days sales in receivables=365/Accounts Receivable Turnover=(365/5.50) 66.36 Days
6) Days Sales in Inventory=365/Inventory turnover ratio
Cost of goods sold=(A) $                                                             15,54,900.00
Average Inventory=(Beginning Inventory+Ending Inventory)/2
Beginning Inventory $                                                               4,52,600.00
Ending Inventory $                                                               5,84,000.00
Average Inventory=(B)=($452600+$584000)/2 $                                                               5,18,300.00
Inventory Turnover=(A)/(B) 3.00 Times
7) Days Sales in Inventory=(365/Inventory Turnover)=(365/3) 121.67 Days
8) Fixed Assets to long term liabilities
Fixed Assets $                                                             36,30,000.00
Long Term Liabilities $                                                             33,00,000.00
(Fixed Assets/Long Term liabilities)=($3630000/$3300000) 1.1
9) Ratio of liabilities to stockholder's equity
Total Liabilities=(A) $                                                             42,82,080.00
Stockholder's Equity=(B) $                                                             53,52,600.00
Total Liabilities/Stockholder's Equity=(A)/(B) 0.80
10) Times Interest Earned=EBIT/Interest
EBIT=(A)/(B) $                                                             11,27,600.00
Interest $                                                               2,64,000.00
Times Interest Earned=(A)/(B) 4.27 Times